OpenStudy (anonymous):

Louise and Molly are twins and their brother Jonathan is four years younger. The total of their three ages is 17 years. (i) Write this information in the form of an equation in 'j', Jonathan's age in years. (ii) What are all their ages?

OpenStudy (anonymous):

6.5

OpenStudy (anonymous):

Sorry?

OpenStudy (anonymous):

Let j be the age of Jonathan J+4 age of twins 2(j+4)+j=17 2j+8+j=17 3j=17-8 3j=9 J=3 Jonathan age is 3 yo, twins are 3+4=7 yo Check: 7+7+3=17

OpenStudy (anonymous):

Thank you! That's really helpful Marina :D x

OpenStudy (anonymous):

You are welcome